Ticket Number: HEA-NOC/19980817-01 Ticket Status: CLOSE
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> UCC
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 19980817 14:55 UTC+1
Ticket Closed: 19980817 15:17:14 UTC Problem End: 19980817 15:47 UTC+1
The line from HEAnet to University College Cork was down.
All traffic to/from UCC.
Informed T.E. (ref. 1364); turned out to be a fibre problem in Cork.
Time to Fix: